# Contest for system observability as an infinitely repeated game

### v2

@@ 5 Illustrative examples
Cases 4-5: let two players start thinking about two-step optimality at different epochs to compare the impact of different players on the equilibrium. Before epoch 20, let two players consider one-step optimality and choose actions according to Algorithm 1 and Algorithm 2. In Case 4, let the attacker consider two-step optimality after 20 epochs and both players consider two-step optimality after 40 epochs. In Case 5, let the defender not the attacker consider two-step optimality after 20 epochs and other settings are the same as Case 4.
Fig. 6 shows the game results for Cases 4-5. Before epoch 20, the game results of both cases are the same as Case 1. From epoch 20-40, the value function evolves in a different loop in Case 4 and the game reaches the Nash equilibrium in Case 5. After epoch 40, the games in both cases reach the Nash equilibrium.
The reason why cases 4,5 reach equilibrium at different epochs is due to the dimension of $\mathcal{V}^{*}$. When the attacker considers two-step optimality at epoch 20, $\operatorname{dim}\mathcal{V}^{*}=3$, which is large and more beneficial to the attacker. In this case, $\operatorname{dim\,}\mathcal{V}^{*}>\min_{H\in\mathbb{R}^{m\times n}}\operatorname{dim\,Ker}\Omega\,(U_{a},H)=1$, thus there is no Nash equilibrium. When the defender considers two-step optimality, $\operatorname{dim}\mathcal{V}^{*}=1$, which is small and more beneficial to the defender. In this case $\operatorname{dim\,}\mathcal{V}^{*}=\min_{H\in\mathbb{R}^{m\times n}}\operatorname{dim\,Ker}\Omega\,(U_{a},H)=1$, thus there is Nash equilibrium. And after epoch 40, there is $\operatorname{dim\,}\mathcal{V}^{*}=\min_{H\in\mathbb{R}^{m\times n}}\operatorname{dim\,Ker}\Omega\,(U_{a},H)=1$ in both cases, thus there is Nash equilibrium.
According to Cases 4-5, the defender can achieve Nash equilibrium by following two-step optimality but the attacker can not, and both cases illustrate the result of Theorem 4.15.
@@
Case 6: let both players consider one-step optimality and choose actions to achieve equilibrium before epoch 40. In epoch 40, when $H=\left[\begin{array}[]{lllll}1&0&0&1&1\\ 0&0&1&0&0\end{array}\right]$, the attacker chooses $U_{a}=\left[\begin{array}[]{lllll}0&1&0&0&0\end{array}\right]\notin BR1^{a}$. Although the value function in epoch 40 reduces, $\operatorname{dim\,}\mathcal{V}^{*}$ in epoch 41 increases, which increases the value function in epoch 42. Thus the equilibrium is broken and the game reaches a new mode: the value function oscillates between 1 and 3. From the Nash equilibrium to the oscillation mode, the average value function increases from 1 to 1.5, which is more beneficial for the attacker. It shows that by giving up the benefit of the current epoch, the attacker can break the Nash equilibrium and make the game result more beneficial for itself. This illustrates the result of Theorem 4.21.
@@ 6 Conclusion
In this paper, we formulate an infinitely repeated game model between the attacker and the defender based on the unobservable subspace dimension. Derivations and algorithms to maximize or minimize the unobservable subspace dimension are given. The game considering one-step optimality and two-step optimality are studied respectively. It is found that the best response sets are not single-valued maps and different choices will lead to different game results. Still, the necessary and sufficient condition for Nash equilibrium is given, which is the same no matter whether players consider one-step optimality or two-step optimality. We find if both players’ actions ensure the optimal of the current epoch, the defender has a dominant role in whether the game can reach an equilibrium. But by giving up the benefit of the current epoch, the attacker can break the Nash equilibrium. Finally, six cases are given to illustrate the main results.

### v3

@@ 5 Illustrative examples
Cases 4-5: We allow two players to apply two-step optimality in different epochs to compare their impact on game result. Before epoch 20, let two players consider one-step optimality and choose actions according to Algorithm 1 and 2. In Case 4, let the attacker consider two-step optimality after 20 epochs and both players consider two-step optimality after 40 epochs. In Case 5, let the defender consider two-step optimality after 20 epochs and other settings are the same as Case 4.
Figure 6  Results of Case 4: the attacker considers two-step optimality after 20 epochs. (a) Evolution of $\operatorname{dim\,Ker}\Omega$; (b) Evolution of $\dim\mathcal{V}^{*}$.
Figure 7  Results of Case 5: the defender considers two-step optimality after 20 epochs. (a) Evolution of $\operatorname{dim\,Ker}\Omega$; (b) Evolution of $\dim\mathcal{V}^{*}$.
Figures 6–7 depict Cases 4–5. Up to epoch 20 both cases follow the trajectory of Case 1. At epoch 20 the attacker (Case 4) or the defender (Case 5) adopts a two-step best response, producing different outcomes. In Case 4, $\dim\mathcal{V}^{*}$ jumps to 3, exceeding $\min_{H}\operatorname{dim\,Ker}\Omega(U_{a},H)=1$; no Nash equilibrium exists and the state cycles until epoch 40. In Case 5, $\dim\mathcal{V}^{*}$ drops to 1 and equals $\min_{H}\operatorname{dim\,Ker}\Omega(U_{a},H)$; the game reaches a Nash equilibrium immediately. After epoch 40 both cases satisfy $\dim\mathcal{V}^{*}=\min_{H}\operatorname{dim\,Ker}\Omega(U_{a},H)=1$ and remain at equilibrium. These results corroborate the necessary and sufficient condition of NE in Theorem 4.7.
Furthermore, in Case 4 the attacker’s two-step move raises the long-run average of $\operatorname{dim\,Ker}\Omega$ from 1.5 to roughly 2.0, whereas in Case 5 the defender’s two-step response lowers it from 1.5 to roughly 1.0. Thus, by selecting an action from its two-step optimal set, either player can shift the long-run value function in its own favor.
@@
Case 6: let both players consider one-step optimality and choose actions to achieve equilibrium before epoch 40. In epoch 40, when $H=\left[\begin{array}[]{lllll}1&0&0&1&1\\ 0&0&1&0&0\end{array}\right]$, the attacker chooses $U_{a}=\left[\begin{array}[]{lllll}0&1&0&0&0\end{array}\right]\notin BR1^{a}$.
Figure 8  Results of Case 6: in epoch 40, the attacker chooses $U_{a}\notin BR1^{a}$. (a) Evolution of $\operatorname{dim\,Ker}\Omega$; (b) Evolution of $\dim\mathcal{V}^{*}$.
The game result of case 6 is shown in Figure 8. Although the value function in epoch 40 reduces, $\operatorname{dim\,}\mathcal{V}^{*}$ in epoch 41 increases, which increases the value function in epoch 42. Thus the equilibrium is broken and the game reaches a new mode: the value function oscillates between 1 and 3, which illustrates the result of Corollary 4.19. From the Nash equilibrium to the oscillation mode, the average value function increases from 1 to 1.5, which is more beneficial for the attacker. It shows that by giving up the benefit of the current epoch, the attacker can break the Nash equilibrium and make the game result more beneficial for itself.
@@ 6 Conclusion
In this paper, we formulate the contest between an attacker and a defender over system observability as an infinitely repeated game whose value function equals the dimension of the unobservable subspace. Derivations and algorithms to maximize or minimize the unobservable subspace dimension are given. Despite the resulting best response sets being set-valued, we established a unified necessary-and-sufficient condition for Nash equilibrium. The long-term outcome of the game depends on whether the two players satisfy the Nash equilibrium conditions. If a defender-dominated Nash equilibrium exists and the defender chooses the corresponding strategy, the controlled invariant subspace collapses immediately to $\{0\}$ and the game enters the lock mode. In this case, the attacker can no longer reduce observability and system security remains permanently fixed at the most favorable level of the defender. Under the more general non-degenerate condition, the game admits two possible outcomes: lock mode and loop mode. Furthermore, we provide a necessary condition for the Nash equilibrium in which the strategies of two players are uncoupled. By deliberately adopting a strategy that violates this condition, the attacker can break the equilibrium. Although the attacker sacrifices an immediate utility, it may achieve a higher long-term value. Finally, numerical case studies confirm these insights.